Spectroscopic observations of compact solar flares

Giulio Del Zanna & H.E. Mason A. Berlicki & B. Schmieder

Department of Applied Mathematics and Theoretical Physics

University of Cambridge Observatoire de Paris, Meudon

Introduction

• The SOHO/Coronal Diagnostic Spectrometer provides essential spectroscopic information to understand the characteristics and dynamics of solar coronal features.

• We give a description of the main observational features and what types of measurements (densities, temperatures, flows) are obtained.

• These provide important observational constraints for flare modeling.

• CDS observations of a compact M1 flare that occurred on October 22 2002, also observed from the ground in chromospheric lines and by RHESSI.

• The flare occurred in a region of strong and mixed magnetic polarity, was compact and lasted only a short time.

Papers – I

- Del Zanna et al. (2004) in preparation

COSPAR 2004: compact solar flares – p.3/22

Papers – II

• On general aspects of quiescent ARs:

-Del Zanna & Mason 2003, A&A, 406, 1089 -Del Zanna 2003, A&A, 406, L5

• On chromospheric evaporation:

-Mason et al. 1986, ApJ, 309,435 -Zarro et al., 1988, ApJ, 333, L99 -Del Zanna et al. (COSPAR 2000), 2002, Adv. Space Res., 30, 551

• Time dependent and non-equilibrium effects: -Bradshaw, Del Zanna & Mason, A&A, 2004, in press

Overview I

Figure 1: THEMIS/MSDP in Na I D_{1} (15:41–16:29 UT) and line-of-sight magnetic field deduced from I and V Stokes overlaid with Hα contours observed by VTT/MSDP at 15:30:01 UT.

COSPAR 2004: compact solar flares – p.5/22

Overview II

• THEMIS observed continuously between 15:41–17:26 UT.

• TRACE and VTT/MSDP observations were performed during the whole flare.

• VTT/MSDP observations in Hα 9 scans between 15:30 and 16:17 UT

• TRACE high cadence (40 s) in the 195 Å band between 15:29 and 15:41 UT

Overview III - GOES

COSPAR 2004: compact solar flares – p.7/22

RHESSI

RHESSI counts rates (the 12–25 keV band are multiplied by a factor of 10).

Movies

MDI MDI movie MDI - during the M1 flare MDI movie TRACE 195 Å TRACE movie

The M1 flare

COSPAR 2004: compact solar flares – p.10/22

The M1 flare

CDS FOV - TRACE 195 Å + contours of Fe XIX intensity

The M1 flare - CDS

COSPAR 2004: compact solar flares – p.12/22

The M1 flare - CDS – Zoom in

COSPAR 2004: compact solar flares – p.13/22

Wavelength calibration

The wavelength calibration was remarkably stable. Uncertainty in the derivation of velocities of the order of 5–10 km/s.

\begin{matrix}λ_{st} (Å) & 520.665 & 522.210 & 542.07 & 550.06 & 553.34 & 555.263 & 557.765 & 558.594 & 562.803 & 592.60 & 624.941 & 629.732 \\ λ (s26055) & 520.675 & 522.208 & 542.031 & 550.070 & 553.365 & 555.252 & 557.758 & 558.590 & 562.791 & 592.590 & 624.936 & 629.735 \\ v (s26055) & 5 & -1 & -21 & 5 & 13 & -6 & -3 & -2 & -6 & -5 & -2 & 1 \\ λ (s26059) & 520.664 & 522.211 & 542.061 & 550.066 & 553.354 & 555.260 & 557.769 & 558.622 & 562.812 & 592.597 & 624.932 & 629.747 \\ v (s26059) & 0 & 0 & -4 & 3 & 7 & -1 & 1 & 15 & 5 & -1 & -4 & 6 \end{matrix}

The Fe XIX case

Our calibration, based on this small sample, provides a wavelength of 592.16 Å, in excellent agreement with Skylab measurements. Lawson & Peackok (1980) reviewed n = 2 → n = 2 transitions, and suggested an energy for the 2s 2 2p 4^{1}D_{2} level of 168770 cm^{−1}. → λ = 592.522 Å → the considerable difference of 183. km/s compared to our result.

Selected areas

Si XII velocity map, with the areas selected for further analysis. Areas A,B,C correspond to regions of strong blue-shifts, while area D of red-shift. Region LP corresponds to the area of peak Fe XIX intensity, while region L1 is where the central part of the flare loop is located. Region BG is a ‘background’ reference region.

COSPAR 2004: compact solar flares – p.16/22 COSPAR 2004: compact solar flares – p.17/22 COSPAR 2004: compact solar flares – p.18/22

Region  A  B C D L1 LP BG
v He I (km/s)  13  -27 -30 -17
v Si XII (km/s) -64  -84 -54 17 -6 -40
v Fe XIX (km/s) -141 -140 -56 22 -13 -26 -

Post-flare CDS raster 15:43–17:07 UT

COSPAR 2004: compact solar flares – p.20/22

Post-flare CDS raster – II

COSPAR 2004: compact solar flares – p.21/22

Summary of preliminary results

• Strong blue-shifts in lines emitted in the 1-8 MK range are observed at the bases of hot Fe XIX structures.

• The hot structures are also clearly visible in the TRACE 195 Å band • The hot structures match the thermal emission as seen by RHESSI • The peak Fe XIX emission, recorded between 15:35 and 15:43 UT,

occurred just above the region of strong mixed polarity.

• The TRACE 195 Å emission is co-spatial with the CDS Fe XIX emission, which suggests that the bulk of the TRACE 195 Å emission at the flare site is due to Fe XXIV.

• Transition region densities can be estimated using O IV.

Acknowledgements Support from PPARC is acknowledged. SOHO is a project of international collaboration between ESA and NASA. We acknowledge the use of the SURF, TRACE on-line databases.